A person puts a few apples into the freezer at -12°C to cool them quickly for guests who are


about to arrive. Initially, the apples are at a uniform temperature of 30°C, and the heat transfer


coefficient on the surfaces is 21 W/m2·°C. Treating the apples as 9-cm-diameter spheres and


taking their properties to be ρ = 900 kg/m3, Cp = 3.81 kJ/kg·°C, k = 0.615 W/m·°C, and α = 1.4×10-7 m2/s, determine the center and surface temperatures of the apples in 1.5 h. Also, determine the amount of heat transfer from each apple. Use Heisler chart to solve the problem.

1
Expert's answer
2022-02-18T15:51:05-0500

Biot number Bi = hLc / k

h = сonvective heat transfer coefficient = 21 W/m2·°C

k = thermal conductivity = 0.615 W/m·°C

LC = characteristic length = volume/area = V/A

For sphere, V = (4/3)πr³ and A = 4πr²

LC = (4/3)πr³ / 4πr² = r/3 = d/6 = 0.09 m / 6 = 0.015 m

Bi = hLc / k = 21 W/m2·°C x 0.015 m / 0.615 W/m·°C = 0.512

Fourier number F0 = αt/LC2 = (1.4×10-7 m2/s x 5400 s) / (0.015 m)2 = 3.36

From the coefficient table used for dimentional approximation of transient 1D heat conduction in sphere it was found "\lambda" = 1.166, A = 1.144

Temperature at the centre of the apple

T(t) = Ae-"\lambda"^2·F (Ti - T) + T = (1.144)e-(1.166)^2(3.36)(30°C + 12°C) - 12°C = -11.5°C

Temperature at the surface of the apple

T(r,t) = Ae-"\lambda"^2·F (Ti - T)(sin"\lambda")/"\lambda" + T = (1.144)e-(1.166)^2(3.36)(30°C + 12°C){sin(1.166)}/(1.166) - 12°C = -11.6°C

Mass of the apple m = "\rho"V

V = (4/3)πr³ = (4/3) x 3.14 x (0.09 m/2)3 = 0.000382 m3

m = "\rho"V = 900 kg/m3 x 0.000382 m3 = 0.344 kg

Maximum heat transfer Qmax = mCp(Ti - T) = 0.344 kg x 3.81 kJ/kg·°C x (30°C + 12°C) = 55.05 kJ

Actual heat transfer of the apple to the surrounding

Q = Qmax [1 - 3 x {(T(r,t) - T) / (Ti - T)} x (sin"\lambda" - "\lambda"cos"\lambda")/"\lambda"3] = 55.05 [1 - 3 x {(-11.6°C + 12°C) / (30°C + 12°C)} x (sin(1.166) - 1.166 x cos(1.166))/(1.166)3] = 54.6 kJ